OpenStudy (anonymous):

Solve the following problems. What is 95% of 280? What is 0.3% of 60? What is 254% of 500? 16 is what percent of 64? 5 is what percent of 950? ( Round to the nearest tenth of a percent) 225 is what percent of 50? 24 is 72% of what number? 3 is 0.02% of what number? 132 is 150% of what number?

OpenStudy (anonymous):

280*.95=266 60*.003=0.18 500*2.54=1270 16/64*100=25% 5/950*100=0.5% 225/50*100=450% 24/.72=33.33 3/.0002=15000 132/1.5=88

OpenStudy (anonymous):

Wow thank you so much! :)